OpenStudy (anonymous):

(x^2 + y^2) dx + (3xy) dy = 0 This equation isn't exact. I am wondering if I should use the equations for a special integrating factor.

OpenStudy (dumbcow):

you can use formulas for bernoulli equations: http://tutorial.math.lamar.edu/Classes/DE/Bernoulli.aspx rearranging gives: \[ \frac{dy}{dx} + \frac{y}{3x} = -\frac{xy^{-1}}{3}\] \[y \frac{dy}{dx} + \frac{y^2}{3x} = -\frac{x}{3}\] make the substitution v = y^2 v' = 2y(dy/dx) \[\rightarrow \frac{v'}{2} + \frac{v}{3x} = -\frac{x}{3}\] now use the integrating factor x^(2/3) \[\rightarrow (x^{2/3} v)' = -\frac{2}{3}x^{5/3}\] integrating \[x^{2/3} v = -\frac{1}{4} x^{8/3}+C\] \[v = -\frac{x^2}{4}+Cx^{-2/3}\] plug "y" back in ..... v=y^2 \[y = \pm \sqrt{Cx^{-2/3}-\frac{x^2}{4}}\]

OpenStudy (anonymous):

Never heard of that and I don't think that is what my prof is looking for

OpenStudy (anonymous):

Calculate the vapor pressure at 20 ∘C of a saturated solution of the nonvolatile solute, urea, CO(NH2)2 in methanol, CH3OH. The solubility is 17 g urea/100 mL methanol. The density of methanol is 0.792 g/mL, and its vapor pressure at 20 ∘C is 95.7 mmHg.